OpenStudy (anonymous):

1 MEDAL HELP! What is one of the solutions to the following system of equations? y2 + x2 = 65 y + x = 7

OpenStudy (haseeb96):

is it y power square?

hero (hero):

@maaryliin, you can only give one medal per solution.

hero (hero):

If others wish to give those who contribute the best response more medals, then they can earn more medals that way. Please do not false advertise medal counts.

OpenStudy (anonymous):

it is y square root of two yes

OpenStudy (shamil98):

y2 + x2 = 65 y + x = 7 lets solve for x. y = -x + 7 (7-x)^2 + x^2 = 65 foil the (7-x)^2 and combine the like terms , you will get a quadratic equation.

OpenStudy (shamil98):

From there use the quadratic formula or factor it ( if possible to factor with a first look ).

OpenStudy (haseeb96):

there are many ways to sloving this type of question in which one is this y^2 + x^2=65 ...........(1) y+x =7 ...............(2) take eq(2) first y=7-x .....(3) put equation (3) in eq(1) (7-x)^2 +x^2 = 65 49 -14x + x^2 +x^2=65 2x^2 -14x =65-49 taking 2 common from left hand side equation 2(x^2-7x)=16 x^2-7x=8 since, 16\2=8 x^2-7x-8=0 now factorized it x^2+x-8x-8=0 x(x+1)-8(x+1)=0 (x-8)(x+1)=0 either, x-8 =0 or x+1=0 x=8 or x=-1 now first we put x = 8 in eq(3) since, y=7-x y=7-8 y=-1 now we put x=-1 in eq(3) since, y= 7-x y=7-(-1) y=7+1 y=8 SOLUTION SETS = {( 8,-1 ) , ( -1,8 ) } this is the correct answer

OpenStudy (anonymous):

yeah but what i dont get is that the problem is multiple choice & this is all it gives me look not the ansewr you got Select one: a. (-8, 1) b. (1, 6) c. (6, 1) d. (9, -2)